OpenStudy (writerforever18):

Factor the following expression. 28u^6v^6-24v^3w^2 Experienced mathematician pls help!!

OpenStudy (anonymous):

have you tried breaking it down

OpenStudy (anonymous):

what do you think it is

OpenStudy (writerforever18):

Well I know that you have to find the GCF between the numbers and then multiply that GCF by the two numbers that multiply together to get the two main numbers in the problem. My concern is what numbers the variables go with.

jhonyy9 (jhonyy9):

28u^6v^6 -24v^3w^2 = 7*4*u^6v^3v^3 -6*4*v^3w^2 = ? - hope these will help you factoring the common terms and so getting the right answer easy good luck bye bye